OpenStudy (abbycross167):

Can someone please help me with this really easy question?? Please? Write the point-slope (y - k = m(x - h)) form of the equation from the information given. Slope = -½; (h, k) = (-1, 2) My Answer~ y-k=m(x-h) m=1/2 h=-1 k=2 How would I put this into an equation?

jimthompson5910 (jim_thompson5910):

is it m = 1/2 or m = -1/2 ? There's a contradiction going on here

OpenStudy (abbycross167):

m=-1/2

jimthompson5910 (jim_thompson5910):

ok so all you need to do really is replace the variables with their set values example: replace m with -1/2

jimthompson5910 (jim_thompson5910):

after you plug in the values, you'll do one simplification

OpenStudy (abbycross167):

so, y-2=-1/2(x-(-1)?

jimthompson5910 (jim_thompson5910):

the `x-(-1)` portion can be re-written into `x+1` since the two negatives cancel to form a positive

jimthompson5910 (jim_thompson5910):

so the answer would be \[\Large y - 2= -\frac{1}{2}(x+1)\]

OpenStudy (abbycross167):

oh ok, is that all?

jimthompson5910 (jim_thompson5910):

yep, that's the point slope form

OpenStudy (abbycross167):

ok great that was really easy!! Thank you sooo very much!!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (abbycross167):

Could you help me with one more problem? I did all the work except one thing